### v1

@@ Abstract
Estimating the difficulty of exam questions is essential for developing good exams, but professors are not always good at this task. We compare various Large Language Model-based methods with three professors in their ability to estimate what percentage of students will give correct answers on True/False exam questions in the areas of Neural Networks and Machine Learning. Our results show that the professors have limited ability to distinguish between easy and difficult questions and that they are outperformed by directly asking Gemini 2.5 to solve this task. Yet, we obtained even better results using uncertainties of the LLMs solving the questions in a supervised learning setting, using only $42$ training samples. We conclude that supervised learning using LLM uncertainty can help professors better estimate the difficulty of exam questions, improving the quality of assessment.
@@ 3.2 Professors’ estimations
Three professors of the University of Groningen were asked to estimate for each question the percentage of students that would answer correctly. All three professors have expertise in Machine Learning and Neural Networks and would be qualified to teach these courses. However, none of them have taught these specific courses, and have never been a student in these courses. This ensures they are fairly knowledgeable about the population of students and the topic, but have not seen students’ performance on these questions. As an example and to provide some calibration, the professors were given one exam question with the true percentage of students that answered it correctly. The professors were also given the correct True/False answers. This was to help them focus on the task — predicting the difficulty of the question rather than solving it. The exact annotation instructions are presented in Table 2.
@@ 4 Results
Method Features NN AML RMSE $\downarrow$ ME $\rho$ $\uparrow$ RMSE $\downarrow$ ME $\rho$ $\uparrow$ Professors Professor 1 0.336 0.048 -0.014 0.357 -0.087 0.062 Professor 2 0.245 -0.038 -0.011 0.185 0.070 0.241 Professor 3 0.184 0.000 0.211 0.193 -0.003 0.224 Average of Professors’ Estimates 0.234 0.001 -0.020 0.205 -0.010 0.173 LLM Direct Estimation Gemini 2.0 Single question 0.242 -0.071 0.118 0.199 -0.061 0.188 All questions 0.218 0.016 0.062 0.307 -0.177 0.052 Gemini 2.5 Single question 0.192 0.035 0.283 0.165 0.044 0.345 All questions 0.199 0.034 0.175 0.243 0.051 0.079 Supervised Learning Dummy Model None 0.174 - - 0.126 - - Linear Regression LLM Uncertainties 0.165 0.024 0.678 0.119 0.004 0.418 TF-IDF 0.174 0.016 0.098 0.126 -0.000 0.251 TF-IDF & Uncertainties 0.165 0.024 0.741 0.119 0.003 0.464 Random Forest LLM Uncertainties 0.170 0.032 0.273 0.120 -0.009 0.282 TF-IDF 0.180 0.004 0.035 0.115 0.003 0.469 TF-IDF & Uncertainties 0.179 0.011 0.049 0.109 0.004 0.464 Support Vector Machine LLM Uncertainties 0.148 0.011 0.811 0.107 0.009 0.582 TF-IDF 0.174 0.018 -0.070 0.117 0.008 0.314 TF-IDF & Uncertainties 0.147 0.001 0.853 0.109 0.013 0.491
@@
Overall, and in line with the study by van de Watering and van der Rijt [7], professors seem to have limited ability to estimate question difficulty. We see that for the Neural Networks (NN) exam, two of the professors estimated $p^{+}\text{-values}$ that do not correlate with the performance of students. Only professor $3$ has a positive rank correlation with $\rho{=}0.211$. This may be partly because professor $3$ did not give an answer to sixteen questions for Neural Networks, presumably the ones they felt uncertain about. For the MSc level Advanced Machine Learning (AML) course professor $2$ and professor $3$ achieved better performances. Their estimated $p^{+}\text{-values}$ consistently show a weak rank correlation with the ground truth.
The Mean Errors are sometimes positive and sometimes negative, depending on the professor and the exam. This suggests that there is no clear pattern of professors consistently over/underestimating question difficulty. The high RMSE does show that overall the professors perform poorly at directly estimating $p^{+}\text{-values}$. Furthermore, for each question we also average the three professor estimates as if they are voting. This did not lead to any improvements.
@@
When assessing the two methods of direct prompting, we find that prompting the model with one question at a time generally leads to lower RMSE and higher rank correlations, with the exception of Gemini 2.0 in the Neural Networks set. Additionally, we find that Gemini 2.5 is consistently more accurate than Gemini 2.0 which corresponds well with their performance in other tasks [18]. When comparing the direct prompting of LLMs to the professors we find that the LLMs tend to perform better. The best LLM method (Gemini 2.5, single question) has better rank correlation than all professors on both exams. For Advanced Machine Learning the best rank correlation from a professor was $\rho{=}0.241$, while that of the LLM was $\rho{=}0.345$.
@@ 5 Discussion
We have shown that Gemini 2.5 is better at question difficulty estimation for Neural Networks and Advanced Machine Learning exams than three university professors. We also find that Gemini 2.5 consistently outperforms Gemini 2.0 on this task, which suggests that future LLM releases may lead to further improvement.
@@
The primary limitation of this study is that we relied on professors that did not teach this specific course. In reality, professors have additional information which can help with the task of question difficulty estimation, such as the performance of a cohort during the semester. At the same time, the professors in our study already have significantly more background information than the better-performing Gemini 2.5, as they are familiar with the rest of the curriculum and know how these students perform in other courses. This limitation may cast doubt whether question difficulty estimation from Gemini 2.5 is better than a professor that has been teaching a specific course. However, it remains clear that the supervised learning method is superior given the large differences.

### v2

@@ Abstract
Estimating the difficulty of exam questions is essential for developing good exams, but professors are not always good at this task. We compare various Large Language Model-based methods with four professors in their ability to estimate what percentage of students will give correct answers on True/False exam questions in the areas of Neural Networks and Machine Learning. Our results show that the professors have limited ability to distinguish between easy and difficult questions and that they are outperformed by directly asking Gemini 2.5 to solve this task. Yet, we obtained even better results using uncertainties of the LLMs solving the questions in a supervised learning setting, using only $42$ training samples. We conclude that supervised learning using LLM uncertainty can help professors better estimate the difficulty of exam questions, improving the quality of assessment.
@@ 3.2 Professors’ estimations
Four professors of the University of Groningen were asked to estimate for each question the percentage of students that would answer correctly. All four professors have expertise in Machine Learning and Neural Networks and would be qualified to teach these courses. One of the four professors (henceforth referred to as Professor 0) was the person who created the exams, has taught the courses and is familiar with the student cohorts that took the exams. The remaining professors have not taught these specific courses, and have never been a student in these courses. This ensures they are fairly knowledgeable about the population of students and the topic, but have not seen students’ performance on these questions, in contrast to Professor 0. As an example and to provide some calibration, the professors were given one exam question with the true percentage of students that answered it correctly. The professors were also given the correct True/False answers. This was to help them focus on the task — predicting the difficulty of the question rather than solving it. The exact annotation instructions are presented in Table 2.
@@ 4 Results
Method Features NN AML RMSE $\downarrow$ ME $\rho$ $\uparrow$ RMSE $\downarrow$ ME $\rho$ $\uparrow$ Professors Professor 0 (Exam Creator) 0.192 0.051 -0.041 0.160 0.038 0.392 Professor 1 0.336 0.048 -0.014 0.357 -0.087 0.062 Professor 2 0.245 -0.038 -0.011 0.185 0.070 0.241 Professor 3 0.184 0.000 0.211 0.193 -0.003 0.224 Average of Professors’ Estimates 0.234 0.001 -0.020 0.205 -0.010 0.173 LLM Direct Estimation Gemini 2.0 Single question 0.242 -0.071 0.118 0.199 -0.061 0.188 All questions 0.218 0.016 0.062 0.307 -0.177 0.052 Gemini 2.5 Single question 0.192 0.035 0.283 0.165 0.044 0.345 All questions 0.199 0.034 0.175 0.243 0.051 0.079 Supervised Learning Dummy Model None 0.174 - - 0.126 - - Linear Regression LLM Uncertainties 0.165 0.024 0.678 0.119 0.004 0.418 TF-IDF 0.174 0.016 0.098 0.126 -0.000 0.251 TF-IDF & Uncertainties 0.165 0.024 0.741 0.119 0.003 0.464 Random Forest LLM Uncertainties 0.170 0.032 0.273 0.120 -0.009 0.282 TF-IDF 0.180 0.004 0.035 0.115 0.003 0.469 TF-IDF & Uncertainties 0.179 0.011 0.049 0.109 0.004 0.464 Support Vector Machine LLM Uncertainties 0.148 0.011 0.811 0.107 0.009 0.582 TF-IDF 0.174 0.018 -0.070 0.117 0.008 0.314 TF-IDF & Uncertainties 0.147 0.001 0.853 0.109 0.013 0.491
@@
Overall, and in line with the study by van de Watering and van der Rijt [7], professors seem to have limited ability to estimate question difficulty. We see that for the Neural Networks (NN) exam, three of the professors (including the exam creator) estimated $p^{+}\text{-values}$ that do not correlate with the performance of students. Only professor $3$ has a positive rank correlation with $\rho{=}0.211$. This may be partly because professor $3$ did not give an answer to sixteen questions for Neural Networks, presumably the ones they felt uncertain about. For the MSc level Advanced Machine Learning (AML) course professor $2$ and professor $3$ achieved better performances, but both underperformed compared to Professor 0, whose estimates moderately correlate with the true $p^{+}\text{-values}$.
The Mean Errors are sometimes positive and sometimes negative, depending on the professor and the exam. This suggests that there is no clear pattern of professors consistently over/underestimating question difficulty. The high RMSE does show that overall the professors perform poorly at directly estimating $p^{+}\text{-values}$. Furthermore, for each question we also average the three professor estimates (excluding Professor 0) as if they are voting. This did not lead to any improvements.
@@
When assessing the two methods of direct prompting, we find that prompting the model with one question at a time generally leads to lower RMSE and higher rank correlations, with the exception of Gemini 2.0 in the Neural Networks set. Additionally, we find that Gemini 2.5 is consistently more accurate than Gemini 2.0 which corresponds well with their performance in other tasks [18]. When comparing the direct prompting of LLMs to the professors we find that the LLMs tend to perform better. The best LLM method (Gemini 2.5, single question) has better rank correlation than all professors on both exams, except when compared to Professor 0, who performs better in the Advanced Machine Learning set.
@@ 5 Discussion
To conclude, in our study we show that strong LLMs estimated exam difficulty at least as well as professors that were not directly involved in the course, while the trained regressor estimated exam difficulty better than the professor who created the exams.
Specifically, we have shown that Gemini 2.5 is better at question difficulty estimation for Neural Networks and Advanced Machine Learning exams than three university professors. We also find that Gemini 2.5 consistently outperforms Gemini 2.0 on this task, which suggests that future LLM releases may lead to further improvement.
@@
The primary limitation of this study is that three of the four professors were not involved in the courses. Naturally, the professor who was actively involved in teaching the courses has additional information which can help with the task of question difficulty estimation, such as knowledge of which subjects are more extensively covered in the lectures. At the same time, all professors in our study already have significantly more background information than the better-performing Gemini 2.5, as they are familiar with the whole curriculum and know how these students perform in other courses. Due to the small scale of this study, its findings cannot be generalized to other educational fields or different types of assessments. Therefore, we want to emphasize that while NLP methods surpassed the performance of some professors in this specific task, it should not be concluded that NLP outperforms all professors in this area.